a 4-letter word is to made using letters from the set (a,b,c,d,e,f) How many words are possible assuming that
a)there are no restrictions
b) no repetition of letters is permitted
c) adjacent letters in the word may not be the same?

Answer by sudhanshu_kmr(1152)


a) no. of possible words when no restrictions = 6*6*6*6 = 1296

b) no. of possible words when no repetition = 6P4 = 360


c) no. of possible words when adjacent letters are not same = 6*5*5*5 = 750